I would to solve this problem that I am stuck on the problem is sequences 2,6,8,14 I need to know the next three numbers in the pattern.

Thank You

Answer by solver91311(24713)

This looks sort of like a variant of the Fibonacci series that starts with 0 and 1 and then each subsequent term is the sum of the two preceeding it. In this case, the series starts with 2 and 6, but it looks like the same rule applies, since 2 + 6 = 8, and 6 + 8 = 14.

So my guess at the next three numbers would be:

8 + 14 = 22
14 + 22 = 38
and
22 + 38 = 60

No guarantees on this one, however.